In a 0.1 M solution of TAPS at 7.9 pH, What fraction of TAPS has its sulfonate group in the R-SO3H form?
Answer #1

pka of TAPS = 8.44

we have buffer pH formula given by Henderson equation

pH = pka + log [conjugate base form] / [ acid form]  

here acid form is RSO3H and its conjugate base form is RSO3-

substituting we get

7.9 = 8.44 + log [ RSO3-] / [RSO3H]

[RSO3-] / [RSO3H] = 10^ ( 7.9-8.44)

[RSO3-] /[RSO3H] = 0.2884

given TAPS Molarity is 0.1 M , hence [RSO3-] + [RSO3H] = 0.1

solving above two equations we get [RSO3H] =0.0776 M and [RSO3-] = 0.0224 M

fraction of RSO3H form = [RSO3H form] / [TAPS] = ( 0.0776 M / 0.1M) = 0.776